### v1

# BLaST: High Performance Inference and Pretraining using BLock Sparse Transformers

Abstract.
The energy consumption of large-scale ML models is dominated by data movement - shuffling billions of parameters across memory hierarchies and data centers. Effective sparsification to prune redundant parameters is still challenging: existing methods incur significant accuracy degradation, performance overhead, or both. We introduce (Bl)ock (a)nd (S)parse (T)ransformers (BLaST), a general, robust, and reliable sparsification method applicable to linear layers in all settings. Our method iteratively sparsifies weight matrices into a block sparsity pattern suitable for efficient sparse matrix-matrix (SpMM) multiplication. BLaST achieves up to 95% sparsity in MLP weights with negligible accuracy loss. Our fused, highly optimized Sparse MLP kernel delivers up to 16.7x speedup over dense MLPs across 9 architectures and 8 datasets, resulting in up to 1.6x inference speedup, 1.11x pretraining speedup and up to 3.12x inference memory usage reduction. BLaST enables the next generation of large-scale AI systems by reducing energy use, memory footprint, and latency.
11footnotetext: Equal contribution
1. Introduction
Figure 1. BLaST can sparsify MLP weight matrices up to 95% with minimal accuracy loss. Our highly optimized SpMM kernel improves end-to-end inference runtime by up to 1.6x of Llama 3.2-1B. Sparsified matrices reduce the number of GPUs required to process Llama 3.2-405B model by up to 2.9x, leading to significant cost reductions.
Figure 2. A high-level overview of the entire BLaST pipeline. Our method offers the flexibility to either leverage pretrained neural network weights for fine-tuning or initialize the model from scratch for pretraining. BLaST performs a forward and backward pass, updating the weights while leveraging a fused, highly optimized Sparse MLP kernel. Then, for MLP blocks BLaST sparsifies the weight matrices with the blocked prune and grow algorithm. First, weights are divided in blocks of size $b\times b$. The blocks with the highest magnitudes, along with those that exhibit the largest gradient magnitudes, are retained. This procedure repeats until the desired sparsity is reached.
Large Language Models (LLMs) based on the Transformer architecture have driven a dramatic expansion in the adoption of AI across everyday applications. State-of-the-art models (Geminiteam2024, ; Touvron2023a, ; Deepseekai2025, ; Radford2018, ; Radford2019, ; Brown2020, ) rely on large-scale data centers for both pretraining on massive datasets, and subsequent deployment to serve millions of users. While recent algorithmic innovations have reduced model size and computational demands (Ainslie2023, ; Meng2025, ), the trend toward ever-larger parameter counts continues unabated (Shen2025, ). Training such models to high accuracy and serving them with low latency requires substantial investment in hardware infrastructure, making cost reduction a central challenge in democratizing AI access.
These hardware investments are overwhelmingly tailored to Transformer-based architectures (Vaswani2023, ). The associated costs scale with the size of the model’s weight matrices: larger models consume more memory, and require more GPUs for storage and computation. At runtime, the dominant energy expenditure stems from data movement—transferring weights within and between memory and compute units on chip and across systems. Revenue generation, in turn, is tightly coupled to throughput, i.e., the number of tokens produced per second. Model pruning (Mishra2021, ; Evci2021, ; Hoefler2021, ; Frantar2023, ; Vanderschueren2022, ; Hubara2021, ; Zhou2021, ; Ashkboos2024, )—removing redundant parameters—is a principled strategy that simultaneously reduces memory footprint, minimizes data movement, and increases inference throughput. As such, it represents a compelling approach for improving efficiency across training, deployment, and serving of LLMs at scale.
A large body of work has shown that dense neural networks contain significant parameter redundancy, and that selectively pruning these parameters can yield smaller models with negligible accuracy degradation (Hoefler2021, ; Koch2007, ; Schotthofer2022, ; Ceruti2022, ; Ceruti2022a, ; Hu2024, ; Hu2024a, ; Hubara2021, ; Zhou2021, ; Zhou2021a, ; Evci2021, ; Frankle2019, ; Child2019, ; Pan2024, ; Mishra2021, ; Ivanov2023, ; Hu2021, ; Buyukakyuz2024, ; Chen2024, ; castro2023venom, ; Frantar2023, ; Ashkboos2024, ; Chen2021a, ; Chen2021, ; Chen2020, ; You2022, ). For instance, Evci et al. (Evci2021, ) prune ResNet (He2015, ) convolution layers by leveraging weight magnitudes and first-order gradients, achieving up to 90% sparsity in pretraining. Sparse tensor operations also require fewer floating-point operations, reducing both compute and data movement.
Despite active research, developing a general-purpose sparsification method that simultaneously delivers high sparsity, high performance, and minimal accuracy loss remains difficult. The primary challenge lies in hardware constraints: modern accelerators rely on dense, batched memory access patterns and computation. For example, Tensor Cores (Markidis2018, ) are optimized for dense matrix fragments, and features like Tensor Memory Accelerators (TMAs) (Nvidia2022, ) require deep memory pipelines to fully utilize available bandwidth. While high sparsity has been demonstrated through unstructured pruning—removing parameters element-wise—such sparsity patterns are difficult to process efficiently on current hardware. Attempts to enforce structured sparsity (e.g., pruning blocks or entire channels) offer better hardware compatibility but typically fail to exceed 25% sparsity without incurring significant accuracy loss (Ashkboos2024, ). The achievable speedup of N:M sparsity (Mishra2021, ; Hu2024a, ; Hu2024, ; castro2023venom, ) is limited due to the application of sparsity only in contiguous blocks of elements in spite of having hardware support for Nvidia GPUs(Mishra2021, ).
To address these challenges, we introduce BLaST, a block sparse neural network framework designed from the ground up with hardware efficiency in mind. The key idea is to jointly optimize for both algorithmic sparsity and hardware execution by integrating aggressive, accuracy-preserving weight pruning with custom high-performance sparse kernels. BLaST identifies optimal pruning strategies while leveraging a specialized SpMM (Sparse Matrix-Matrix Multiplication) kernel that outperforms even vendor optimized dense (NVIDIA_cuBLAS_2025, ) and sparse (Abdelfattah2024, ; cusparse, ; hipsparse, ) implementations, delivering significant gains. Our kernel accelerates matrix multiplication by up to 16.7x, enhances the MLP layers in Llama-405B by 8.8x, and boosts inference performance by 1.6x—thus enabling the development of trillion-parameter models. Figure 1 shows the key improvements by BLaST to the Llama3.2(Touvron2023a, ) family of models. The high level design of BLaST is presented in Figure 2.
BLaST provides a unified solution for efficient machine learning across pretraining, fine tuning, and post-training compression. We integrate sparsity into the entire model lifecycle while aligning with hardware execution constraints. During pretraining, BLaST applies a custom, hardware-friendly prune-and-grow strategy that incrementally removes redundant weights, reaching up to 80% sparsity with minimal performance overhead—each pruning step simultaneously improves the efficiency of forward and backward passes. In the fine tuning stage, BLaST builds on the sparsified model, and integrates seamlessly with existing training pipelines [^1] https://huggingface.co/docs/transformers/training, achieving up to 95% sparsity while increasing inference throughput. For post-training model compression, BLaST employs knowledge distillation to generate a compact surrogate model from a dense pretrained one, preserving accuracy while reducing memory footprint, and improving forward pass latency the same as in the fine tuning stage.
The custom SpMM kernel developed in BLaST is also available as a standalone sparse linear algebra primitive, designed for use beyond AI workloads. Here, we exploit block sparsity to unlock unprecedented efficiency across a broad range of scientific and industrial applications where sparse computation is increasingly critical. BLaST’s blocked SpMM (BSpMM) sets a new benchmark, outperforming all existing SpMM libraries—both open-source and commercial. Remarkably, for favorable input structures, BSpMM can even surpass CUTLASS in dense matrix-matrix multiplication performance. On average, as little as 50% sparsity is sufficient to outperform vendor-optimized GEMM libraries, and at 95% sparsity—commonly achieved by BLaST—BSpMM delivers up to 16.7x speedup over the best existing dense and sparse alternatives.
In summary, this paper makes the following contributions:
  1. (1)
    We introduce BLaST, a block sparse neural network framework designed from the ground up for hardware efficiency. BLaST provides a unified, end-to-end solution for efficient machine learning, covering the full model lifecycle—including pretraining, fine tuning, and post-training compression—while maintaining high accuracy under extreme sparsity.
  2. (2)
    We develop a high-performance, scalable implementation of BLaST, including a custom block-sparse SpMM kernel optimized for modern GPU architectures that significantly outperforms both dense and existing sparse baselines.
  3. (3)
    We conduct an extensive empirical evaluation across 15 model architectures and 9 datasets, demonstrating that BLaST achieves up to 90–95% sparsity with no accuracy degradation. Our results show up to 16.7x speedup in kernel performance and up to 1.6x improvement in end-to-end training and inference throughput. BLaST achieves up to 3.12x reduction in inference memory footprint, resulting in 2.9x reduction in GPU setup and operating costs.
2. Background
We now review the key components of the Transformer-based models. We also analyze them from the scope of sparsification strategies and lay foundations for the BLaST approach.
2.1. Transformer-based ML models
The Transformer architecture (Vaswani2023, ) has emerged as the foundational building block of state-of-the-art large language models (Touvron2023a, ; Deepseekai2025, ; Geminiteam2024, ),vision models (dosovitskiy2020image, ), models for science (jumper2021highly, ; sonderby2020metnet, ), and multimodal models
 (radford2021learning, ; ramesh2021zero, ). Each layer of the Transformer architecture comprises two main components: Attention and MLP, plus normalization layers (Figure 2).
Attention. While there are many variants of the Attention module: multi-head attention (Vaswani2023, ), multi-query attention (shazeer2019fast, ), grouped-query attention (ainslie2023gqa, ) and multi-latent attention (liu2024deepseek, ), its underlying compute pattern is the multiplication of transient tensors $Q$, $K$, and $V$ that are computed based on the activation tensor $X$ and trainable weight matrices. The Attention layer captures the relations between tokens - the output state of the token after the attention module depends on the values of other tokens.
The MLP layer, on the other hand, multiplies the activation tensor $X$ directly with weight matrices. Depending on the architecture, typically there are two (radford2019language, ) or three (Touvron2023a, ) weight matrices. Arguably, MoE (mixture-of-experts) (shazeer2017outrageously, ) can be viewed as a functional equivariant of MLP: the essence being that each token is processed independently - the state of the token after MLP depends only on its previous state and the weight matrices, but does not depend on other tokens. For the Llama-family models, the MLP operator is defined as:
(1) $Y=\text{MLP}(X)=\left(SiLU\left(XW_{1}\right)\odot\left(XW_{2}\right)\right)W_{3}$
where $SiLU$ represents the Sigmoid Linear Unit activation function, $X$ represents the input to the MLP layer, and $W_{1}$, $W_{2}$, and $W_{3}$ are the weight matrices. The MLP layer accepts a tensor $X,X\in\mathbb{R}^{n\times s\times e}$ of size $n\times s\times e$, where $n,s,e$ are batch size, sequence length, and embedding size, respectively. Then the model performs a series of expansions and contractions on this input $X$ to generate the output $Y$ as shown in Equation 1.
2.2. Model Sparsification
Related works on sparsification in deep learning suggest that no single method is universally applicable across all scenarios (hoefler2021sparsity, ). Various techniques have been explored—from unstructured weight pruning (Frankle2019, ; prasanna2020bert, ; gordon2020compressing, ; Chen2020, ; guo1909reweighted, ; sanh2020movement, ), which eliminates redundant individual connections, to structured approaches that remove entire neurons (polyak2015channel, ; hu1607network, ; pan2016dropneuron, ; lee2018snip, ), filters (luo2020autopruner, ; luo2017thinet, ; he2019filter, ; you2019gate, ), or methods that sparsify attention (michel2019sixteen, ; xu2025xattention, ; voita2019analyzing, ; mccarley2020structured, ; yun2020n, ; parmar2018image, ; guo2019star, ; child2019generating, ; beltagy2020longformer, ; zaheer2020big, ; li2020sac, ; zhao2019explicit, ). Recently, XAttention (xu2025xattention, ) has proposed block sparse attention with antidiagonal scoring, and NSA (yuan2025native, ) introduces a natively trainable sparse attention mechanism – both for long-context Transformers.
However, empirical studies indicate that either sparsifying MLP weights or adopting a Mixture-of-Experts (MoE) tends to be more robust, leading to substantial reductions in memory footprint and improvements in parallelization (shazeer2017outrageously, ; lepikhin2020gshard, ; fedus2022switch, ; huang2024toward, ). In contrast, approaches focused on sparsifying attention components mainly reduce the size of transient storage during computation without significantly lowering the overall parameter count. These observations underscore the need for flexible, multi-strategy sparsification approaches that balance efficiency gains with the preservation of network trainability and generalization.
In this work, we propose a general solution that is applicable across the entire range of current and future Transformer architectures. Our proposed solution remains accurate as sparsity in the MLP layers increases, thereby ensuring scalable performance for the future of AI.
3. BLaST - (BL)ock (S)parse (T)ransformers
We now proceed to describe BLaST — a novel method designed to improve both inference and pretraining execution time, while also reducing the memory footprint. Section 3.2 introduces the algorithm to obtain the neural networks in the desired block sparse pattern. Section 3.3 introduces our optimized BSpMM kernel.
3.1. Overview
Figure 2 illustrates the high-level pipeline of BLaST for pretraining and finetuning. At the start of training, if pretrained weights are available, they are used to initialize the Transformer-based architecture. Next, BLaST performs a forward and backward pass using our optimized BSpMM kernel (Section 3.3), updating the parameters. Then, the blocked prune and grow algorithm (Section 3.2) is applied in the MLP blocks, adaptively pruning and regrowing weights. This process repeats until the desired sparsity level is reached.
3.2. Blocked Prune and Grow
The “early bird” lottery ticket hypothesis (Chen2021, ; You2022, ) proposes that certain weights are more important than others, and that they can be selected during pretraining without significantly affecting the model’s accuracy. According to this hypothesis, winning weights are chosen by pruning the weights with the smallest magnitudes (Frankle2019, ; Chen2021a, ; Chen2021, ; You2022, ). RigL (Evci2021, ; Evci2022, ) demonstrated that the pruned model’s accuracy can be further improved by incorporating gradient flow information to select the winning weights. This improvement is achieved by applying the same pruning function, $S()$, to both the weights and their corresponding gradients. In the final step, a weight is retained if it meets one of two conditions: it is either directly preserved by the function $S()$ or regrown based on gradient information.
Using RigL on CNNs (Evci2021, ; Evci2022, ) results in fine-grained unstructured sparsity of the weight matrices. Although this results in a reduced number of floating point operations during the forward and backward propagation, it does not translate into improved performance due to memory bandwidth limitations of modern GPUs(Abdelfattah2024, ; naumov2010cusparse, ; hipsparse, ). As a result, we propose to prune the weight matrices of the MLP layer of LLMs using block sparse matrices as shown in the bottom part of Figure 2.
Our blocked prune and grow algorithm gradually reduces the number of parameters of the weight matrix $W$ until a sparsity $s_{max}$ has been achieved over $m$ iterations. It can be implemented as a simple extension to a forward and backward training loop. Listing LABEL:lst:loop-prune-and-grow shows the blocked prune and grow procedure with a regular training loop as Python code. The training loop and optimizer are executed in forward_and_backward_step(). The masks are updated once every $step\_size$ iterations. The masks for each weight matrix are stored in boolean tensors and applied to the weight matrices in the prune_weights() step.
1for iteration in range(train_iters):
2 forward_and_backward_step()
3 if iteration % step_size == 0:
4 generate_masks()
5 prune_weights()
Listing 1: Training loop of our blocked prune and grow algorithm. The actual grow and prune procedure is implemented in the function generate_masks().
Figure 2 illustrates the prune and grow procedure implemented by the function generate_masks() for a weight matrix $W_{i}$ and its corresponding gradient $G_{i}$ for a single iteration. First, the matrices are interpreted as blocks of size $b\times b$, after which the pruning function $S()$ is applied to both $W_{i}$ and $G_{i}$. The function $S()$ computes the Frobenius norm of each block and drops the blocks with the least norm in order to maintain a sparsity $s_{i}$ as given by Equation 2. The pruned blocks which will be zeroed are represented by the white blocks in Figure 2. The third difference step computes the set difference between the non-zero block indices in the $S(W_{i})$ and $S(G_{i})$ matrices, denoted by the red blocks in matrix $D_{i}$. These blocks are regrown back into the sparse weight matrix $S(W_{i})$ in the fourth step, which results in the newly pruned weight matrix $W_{new}$. The regrown blocks added to $W_{new}$ are initially set to zero and then updated during subsequent optimization steps in training. The initialization to zero is done so that the regrown weights do not contribute to the transformation applied by the matrix. Note that in our actual implementation we use boolean masks throughout the pruning and growing process to save time and memory. The $W_{new}$ matrix is ultimately generated in the prune_weights() function.
The weight sparsity $s_{i}$ at iteration $i$ is determined by Equation 2 (Zhu2017, ).
(2) $s_{i}=s(i)=s_{max}+(s_{init}-s_{max})\left(1-\frac{i}{m-d}\right)^{3},$
where $s_{init}$ represents the initial sparsity, which determines the sparsity at the beginning of the training, $m$ is the maximum number of training iterations, and $d$ is a decay term that controls how fast the maximum sparsity $s_{max}$ is reached.
Unlike pruning algorithms that use straight-through estimators (Hu2024a, ; Hubara2021, ; Bengio2013, ), our method prunes the weight matrices directly. Consequently, the same pruned weight matrix is applied during both forward and backward propagation. During training, the dense weight and gradient matrices are kept intact, while the pruned weight $W_{new}$ is stored in the blocked compressed sparse column (BCSC) format [^2] https://pytorch.org/docs/stable/sparse.html#sparse-bsc-tensor for forward and backward propagation. Although this leads to higher memory consumption during training (which gradually decreases as the sparsity $s_{i}$ increases), it ultimately yields improved speedup and reduced inference memory usage, as shown in Section 5.
3.3. BSpMM Kernel
We now introduce our custom BSpMM kernel - the elementary algebraic building block that secures maximal speedup over dense baselines. The BSpMM kernel, while being part of our BLaST framework, can also be used independently as a stand-alone SpMM kernel, paving the road for fast, sparse linear algebra kernels across various domains. We then further extend it to the MLP blocks inside Transformer architectures to capture the most common compute pattern in ML workloads.
3.3.1. Kernel Outline
Figure 3. Overview of BLaST BSpMM. Our kernel employs a blocked Compressed Sparse Column (CSC) format, obtained by sparsifying neural network weights. BLaST leverages block-level, bottom-up 2D parallelism—implemented in Triton–to maximize GPU hardware resource utilization.
Different internal storage formats are preferred depending on whether the sparse matrix $W$ is multiplied from the left ($Y=XW$) or from the right ($Y=WX$). Since the PyTorch implementation of a torch.nn.Linear linear layer assumes multiplying from the left, for brevity, we explain only this variant. The $Y=WX$ kernel is symmetrical, so we omit it from this discussion.
Figure 3 shows the high-level design of our BLaST BSpMM kernel based on the SMaT framework (okanovic2024high, ). SMaT is an SpMM library that exploits Tensor Cores (TCs) for the efficient multiplication of unstructured sparse matrices. The block‑sparse implementation of SMaT relies on the low‑level CUDA Matrix Multiply‑Accumulate (MMA) API to extract maximum performance from GPUs. Its superior performance is attained by efficiently streaming thread‑level blocks, overlapping computation with data transfers, and fully engaging all available TC units through bottom‑up two‑dimensional parallelism (kwasniewski2019red, ) .
The main difference between BLaST BSpMM and SMaT BSpMM is that the sparsity structure is not imposed by the unstructured CSR input but rather arises from our custom sparsification strategy. As a result, we can bypass the key performance bottlenecks identified by Okanovic et al. (okanovic2024high, )—namely, the excessive fill-in caused by imposing an explicit block structure on CSR, and the highly skewed load imbalance among warps. The design of BLaST’s BSpMM kernel is guided by a rigorous performance model (okanovic2024high, ).
The carefully tuned rectangular block shape results from optimizing shared memory and register usage, while minimizing accuracy loss due to overly aggressive sparsification. The two block dimensions—blk_M, and blk_N—are optimized for compute performance, and model accuracy, respectively. Note that blk_N corresponds to block parameter $b$ introduced in Section 3.2.
  1. (1)
    blk_N determines the sparse block size. When the block is smaller, the sparsification strategy has more freedom to choose the best blocks to prune—resulting in a more fine-grained masking process. Conversely, larger blocks enable more efficient computation. We empirically select the largest block sizes that preserve high accuracy. This will be further explored in Section 5.4.1.
  2. (2)
    blk_M controls the number of rows from the dense matrix that participate in the threadblock-level reduction, thereby increasing the reuse of data loaded into shared memory from the sparse block. We use the COSMA analytical compute model (kwasniewski2019red, ) to derive the optimal block size, given hardware constraints such as the number of available registers, shared memory capacity, and warps per threadblock.
3.3.2. Implementation
To secure architecture-agnostic implementation and kernel portability, BSpMM is implemented in Triton (tillet2019triton, ). Triton is well-suited for block computations, using advanced heuristics to automatically generate efficient accelerated code for both NVIDIA and AMD GPUs. It seamlessly utilizes all modern features of the target hardware, such as asynchronous pipeline loads, Tensor Memory Accelerators (TMAs) (Nvidia2022, ), or Tensor Core (Markidis2018, ) computations. Furthermore, Triton rearranges threads within a thread block (called program instance) to avoid bank conflicts, vectorize memory accesses, and align registers for Tensor Core’s fragments.
The main challenge to implement a high-performance BCSC-based, batched SpMM kernel is to handle inherent data indirection efficiently. Compared to dense $Y=XW$, all memory accesses are known at the compile time, and we can create a provably optimal static schedule with compile time tilings, data distribution, and memory pipelining (ziogas2022deinsum, ). The sparse data structure prevents the static access pattern of the dense matrix $X$, since, to achieve data movement optimality, only necessary blocks of $X$ can be loaded - and the sparsity structure of $W$ determines this. We implement this in Triton by determining the dynamic range of accesses to $X$ with pointer algebra on the blk_col_ptr pointer array. The main part of the kernel is presented in Listing LABEL:lst:spmm-kernel. The Triton API can be found in its documentation (tillet2019triton, ).
1for k in range(num_blocks):
2 # load BCSC block
3 W_block = tl.load(block_vals_ptrs + k * blk_stride)
4 # calculate the block row of the current BCSC block
5 row_idx = tl.load(blk_row_idx + k)
6 # move X_ptrs to the correct block
7 X_ptrs = X_ptrs_base + row_idx * blk_K * stride_xk
8 # load dense batched matrix (activation tensor X)
9 X_block = tl.load(X_ptrs)
10 accumulator = tl.dot(X_block, W_block, accumulator)
11 # Advance the ptrs to the next K block.
12 blk_vals += blk_M * blk_stride
Listing 2: Main compute loop of the BSpMM kernel
3.3.3. Block Sparse MLP
In the MLP block, the linear layer is followed by a nonlinear operator. In most of machine learning, nonlinear operators fall into two categories – element-wise nonlinearities, such as ReLU, GeLU, SiLU, and row-wise operators, that perform some type of reduction, usually over the embedding dimension. Examples of such nonlinear operators are LayerNorm (Ba2016, ) or RMSNorm (Zhang2019, ).
While tensor contractions are usually compute-bound (even in the sparse case), nonlinear operators are memory-bound. The arithmetic intensity — the ratio of arithmetic operations per loaded element — is a small constant, usually 1 or 2. Therefore, to prevent a memory bottleneck, nonlinear operators have to be fused with the compute-bound operations.
Our optimized BSpMM kernel for MLPs fuses the nonlinearities that follow the SpMM operator into one kernel. Our implementation supports all common nonlinearities, further improving performance and reducing data movement.
4. Evaluation
We conduct a comprehensive set of experiments to evaluate the performance and accuracy of BLaST. Our experiments span 8 different datasets in both text and vision domain, across 9 different Transformer architectures.
In Section 5, we train the models on a diverse set of benchmark datasets, including OpenWebText (Gokaslan2019OpenWeb, ), CIFAR-10 (krizhevsky2009learning, ), LAMBADA (paperno2016lambadadatasetwordprediction, ), and five tasks from the GLUE benchmark (wang2019gluemultitaskbenchmarkanalysis, ). Furthermore, we conduct experiments using the following Transformer architectures along with their respective parameter counts: Llama3.2 (Panferov2025, ) (1.59B), GPT2-small (Radford2019, ) (124M), GPT2-medium (Radford2019, ) (355M), GPT2-large (Radford2019, ) (774M), GPT2-XL (Radford2019, ) (1.44B), ViT-B/16 (dosovitskiy2020image, ) (86M), ViT-B/32 (dosovitskiy2020image, ) (86M), ViT-L/16 (dosovitskiy2020image, ) (307M), ViT-L/32 (dosovitskiy2020image, ) (307M). In all our experiments, we adhere to the scientific benchmarking guidelines proposed by Hoefler et al. (hoefler2015scientific, ). That is, when reporting speedups we report the single parallel process as the base case.
Hardware Infrastructure.
We evaluate BLaST on a compute partition of the Alps supercomputer [^3] https://www.cscs.ch/computers/alps equipped with 2,688 compute nodes. Each node is equipped with four NVIDIA Grace Hopper GH200 superchips interconnected with HPE Cray Slinghot-11.
Software Stack.
All experiments were executed using Pytorch v2.7.0, CUDA v12.8, Megatron-core v0.10.0, Apex v0.1 and Transformers_engine v2.0.0.
5. Results
In Section 5.1, we present the speedups achieved by BLaST’s BSpMM compared to state-of-the-art sparse matrix-matrix multiplication algorithms, as well as the improvements observed within the MLP layers of LLMs scaled to 405B parameters. This section also showcases the end-to-end speedup achieved during LLM inference. Section 5.3 demonstrates the accuracies achieved during pretraining using BLaST’s Blocked and Prune algorithm. In Section 5.2, we show that BLaST can be used to recover the accuracy of previously pretrained models when inference speedup and memory reductions are of interest. Lastly, Section 5.4 presents an ablation study of BLaST, demonstrating its robustness to the choice of hyperparameters.
5.1. BLaST BSpMM Speedup
We begin by analyzing the speedup achieved by the BLaST BSpMM kernel relative to state-of-the-art matrix multiplication algorithms. Throughout this section, we report the speedup as the ratio
$\frac{\min\left(t_{CUTLASS},\,t_{cuBLAS}\right)}{t_{BLaST{}}},$
where $t_{CUTLASS}$ represents the execution time using the CUTLASS library, $t_{cuBLAS}$ denotes the time for cuBLAS, and $t_{BLaST{}}$ is the execution time for our method. In all experiments, we present the results for blk_M = blk_N $\in\{32,64,128\}$ which coincides to the TC MMA input sizes (see Section 3.3) as the sparsity level increases.
Figure 4 shows the speedup of the BSpMM kernel when multiplying individual matrices corresponding to typical embedding and context length sizes in LLMs. In a matrix multiplication routine characterized by the dimensions $M$, $N$, and $K$, the $Emb$ dimension corresponds to $K$, $Seq$ corresponds to $M$, and $N$ is set to $4\times Emb$. This indicates that as both the $Emb$ and $Seq$ dimensions increase, the speedup also increases. BLaST BSpMM is faster by as much as 16.7x for $Emb=32,768$, $Seq=1,024$, and sparsity 95%. We observe that, in general, higher performance is achieved with a larger block size; this plot uses a 128×128 block size. Moreover, we observe that BLaST SpMM with a 128×128 block size can even outperform cuBLAS by 1.1x in the fully dense case (0% sparsity). Note that BLaST BSpMM supports all data types supported by Triton (tillet2019triton, ). Due to the space constraints, we include the results only for the most popular data type in ML workloads: BF16. Using lower precision, e.g. torch.float8_e5m2, we measured up to 23x speedup for $Emb$ = 32,768, $Seq$ = 1,024, and sparsity 95%.
Lastly, we evaluate the speedup of BLaST BSpMM in the Llama model family, ranging from the Llama-3.2 1B to the Llama-3.1 405B. Figure 5 demonstrates the speedup in the MLP blocks. We observe that even 70% sparsity in the MLP blocks is sufficient to achieve a 2x speedup. As the models scale up to 405B parameters, we achieve up to an 8.8x speedup. In Figure 6, we show the end-to-end speedup in the inference of the Llama-3.2 1B model across different block sizes. One can achieve up to a 1.6x speedup even for a smaller model such as the Llama-3.2 1B using our proposed BSpMM. In conclusion, these experiments demonstrate that BLaST represents a promising step towards scaling models to the trillion-parameter era.
Figure 4. Ablation study on the BSpMM kernel speedup for various block sizes and embedding sizes, showing scaling up results for LLMs.
Figure 5. MLP layer speedups for the Llama family with BLaST 128x128, reaching 8.8× for Llama-3.1 405B
Figure 6. Inference speedup for Llama-3.2 1B for various block sizes and sparsity levels. For 70% sparsity, BLaST yields 1.3x speedup, and for 95% sparsity, BLaST reaches up to 1.6x speedup.
5.2. BLaST for Inference
Pretraining neural networks from scratch demands substantial computational resources, especially for large dense models. Given the widespread availability of pretrained LLMs—which embody significant compute investments—there is a strong incentive to leverage these models rather than retrain a network from scratch. Our approach is capable of knowledge transfer by initializing the network with weights from an existing pretrained dense model. These pretrained weights may originate either from a dense network trained on the same dataset (the post-training compression setting) or from a network pretrained on a different dataset (the fine tuning setting). When training with pretrained dense weights, we have observed that, in cases involving large vocabulary sizes, utilizing a knowledge distillation loss improves the final accuracy (muralidharan2024compact, ; blakeman2025nemotron, ). That is, rather than using solely the cross-entropy loss $L_{CE}$, we backpropagate using a combined loss $\alpha L_{CE}+\beta L_{KL}$, where $L_{KL}$ represents the Kullback-Leibler divergence computed between the BLaST logits and those of the dense pretrained model, with weighting coefficients $\alpha,\beta\in\mathbb{R}$. In our experiments, we do not conduct an ablation study on the effect of knowledge distillation, as it is beyond the scope of this paper; instead, we report the best performance achieved between using, and not using the knowledge distillation loss.
In all experiments, we iteratively sparsify the model while simultaneously training it to recover the accuracy of the dense version (see Section 3). BLaST yields a model that maintains high accuracy while benefiting from faster inference and a reduced memory footprint. We demonstrate the capability of BLaST to recover the full accuracy of a pretrained dense network on downstream tasks.
In Table 1 we compare the performance of sparsifying BLaST under different levels of sparsity, and various block sizes on a subset of the GLUE benchmark.
BLaST CoLA SST-2 MRPC RTE WNLI Avg.
$s_{max}$ $b$ Matt. Corr ACC ACC / F1 ACC ACC Score
70% 32 41.59 84.98 72.06/80.94 59.57 56.34 65.91
80% 32 39.21 85.89 68.14/77.35 61.01 56.34 64.66
90% 32 27.78 84.63 70.34/79.17 61.37 56.34 63.27
95% 32 28.87 85.32 70.83/79.73 59.21 56.34 63.38
70% 64 34.22 84.40 66.42/77.28 59.21 56.34 62.98
80% 64 32.81 85.67 69.36/79.13 60.29 56.34 63.93
90% 64 24.01 84.29 68.87/77.36 58.12 56.34 61.50
95% 64 30.53 85.32 68.38/78.96 55.23 56.34 62.46
70% 128 36.70 86.24 70.83/79.86 60.65 56.34 65.10
80% 128 34.80 85.55 70.59/79.87 53.07 56.34 63.37
90% 128 34.76 85.09 69.12/80.25 59.21 56.34 64.13
95% 128 31.22 83.94 68.87/79.81 54.87 56.34 62.51
Dense 36.44 87.04 73.28/81.22 62.45 56.34 66.13
Table 1. Fine tuning Llama 3.2 1B with BLaST for GLUE benchmark. BLaST is robust to the choice of tile size, capable of lossless sparsification up to 95% for WNLI, and even outperforming Dense for CoLA.
The GLUE benchmark assesses the general language understanding capabilities of models. Specifically, we evaluate on five tasks—Recognizing Textual Entailment (RTE), the Multi-Genre Natural Language Inference Corpus (MNLI), The Stanford Sentiment Treebank (SST-2), the Winograd Schema Challenge (WNLI), and the Corpus of Linguistic Acceptability (CoLA)—and report the average score calculated as the arithmetic mean of all other task specific metrics. We observe that BLaST is robust to variations in block size, achieving a similar average score across the five tasks regardless of the chosen block size. Furthermore, BLaST maintains an accuracy comparable to that of the dense baseline. For instance, on the WNLI dataset, BLaST achieves performance on par with the dense model across all settings, and on the CoLA dataset, the variants BLaST-70%/32x32, BLaST-80%/32x32, and BLaST-70%/128x128 outperform the dense baseline.
Collectively, these results demonstrate the robustness of BLaST to varying sparsity levels, and block sizes for reconstructing accuracy, making BLaST a ready to use method for pretrained networks when speedup and memory reduction are of importance.
5.2.1. Memory Footprint for Inference
The pruning of the weights by up to 95% can have a large impact on the total number of parameters of the model, and with that affect model’s memory footprint. Figure 7 shows the number of NVIDIA GH200 GPUs that are necessary for storing the model weights, assuming 96GB HBM per GPU and FP32 weights. BLaST is capable of reducing the number of necessary GPUs to almost a third of the dense Llama 3.2 405B. This directly translates into reduced capital investments required for establishing inference data centers and lower energy consumption for their operation.
Figure 7. Number of GH200 GPUs required for storing weights in FP32 of the given model assuming 96GB per GPU.
5.3. BLaST for Pretraining
(a) GPT2-XL pretraining with 4.9B tokens.
(b) Llama 3.2 pretraining with 4.2B tokens.
Figure 8. Time per iteration for GPT2-XL and Llama 3.2 with OpenWebText (Gokaslan2019, ). Each line plot uses the hyper parameters from Table 2.
5.3.1. Pretraining LLMs
We pretrain GPT2 and Llama 3.2 on the OpenWebText (Gokaslan2019, ) dataset. The dataset consists of 8,012,769 documents (9,035,582,198 tokens), out of which 4,007 documents are used for building the test dataset. All GPT2 models are trained on 4,915,200,000 tokens (4.9B) and Llama 3.2 is trained on 4,194,304,000 (4.2B) tokens. The test dataset is the same for all models. Section 5.3.2 shows the performance and accuracy for pretraining GPT2 and Llama 3.2 models. The parameters used for controlling the sparsity are found in the ablation studies in Section 5.4.
5.3.2. Accuracy and Performance of Pretraining
Table 2 shows the end-to-end time and perplexity for GPT2-XL and Llama 3.2 models with BLaST configurations. The GPT2-XL model pretraining shows a maximum speedup of 10.3% using 80% maximum sparsity (②) compared to the dense model (①).
{tblr}
cells = c, cell21 = r=3, cell29 = r=3, cell210 = r=3, cell51 = r=2, cell59 = r=2, cell510 = r=2, hline1-2,5,7 = -, Model & $b$ $s_{max}$ $step\_size$ $d$ $L$ BLaST
BLaST
PPL
Dense
Dense
PPL
GPT2-XL 128 80% 100 9000 2 49.36 ② 5.19 54.45 ① 4.79
128 75% 100 9000 2 49.86 ③ 5.14
64 70% 100 0 2 53.00 ④ 5.11
Llama3.2 128 80% 25 200 2 68.79 ⑥ 59.74 70.83 ⑤ 54.70
128 55% 50 200 2 70.33 ⑦ 58.2
Table 2. Pretraining results for BLaST with GPT2-XL and Llama 3.2 on OpenWebText. BLaST reduces end-to-end time of pretraining while preserving accuracy.
Figure 8(a) shows the time per iteration over 4.9 billion tokens for the GPT2-XL model. The performance is compared for various sparsity configurations of BLaST against the dense case. The spikes observed for the sparse configurations are caused by mask generation, which occurs every 100 iterations. Dense matrix multiplication is used until the sparsity reaches 60%. At 60% sparsity, BSpMM routines (Section 3.3) become active, which explains the rapid decrease in per-iteration time as the sparsity increases. BLaST-80%/128x128 (plot ②) and BLaST-75%/128x128 (plot ③) configurations exhibit an earlier activation of SpMM routines due to higher values of $d$. Figure 8(b) shows the pretraining performance of Llama 3.2 over 4.2 billion tokens.
In Table 2, $s_{max}$ represents the maximum permissible sparsity in a given linear layer (see Equation 2). The sparsity-controlling hyperparameters include $b$ (the block size - Section 5.4.1), $step\_size$ (Section 5.4.2), $d$ (the sparsity decay - Section 5.4.3), and $L$ (the number of dense layers - Section 5.4.4).
5.3.3. Pretraining ViTs
In addition, we demonstrate that BLaST can be applied not only to NLP tasks but also to the vision domain. We train a ViT (dosovitskiy2020image, ) on CIFAR-10 (krizhevsky2009learning, ). Table 3 shows a comparison of four different ViT architectures under increasing sparsity levels achieved by BLaST.
Model Dense BLaST-70% BLaST-80% BLaST-90% BLaST-95%
ViT-B/16 98.57% 93.45% 93.82% 94.56% 93.08%
ViT-B/32 98.54% 92.91% 92.95% 92.70% 91.49%
ViT-L/16 98.88% 95.82% 95.71% 95.48% 95.31%
ViT-L/32 98.75% 92.13% 93.36% 93.59% 93.57%
Table 3. Accuracy of training four ViT architectures with BLaST. ViT is pretrained on a large vision task (dosovitskiy2020image, ), demonstrating BLaST’s knowledge transfer capability.
These results demonstrate the robustness of BLaST in handling increased sparsity levels in the MLP blocks of the ViT architecture.
Furthermore, in Figure 9 we illustrate how test accuracy evolves for both BLaST and the dense network as we iteratively train and sparsify the ViT architecture on CIFAR-10. We plot the test accuracy after every epoch (i.e., after each complete pass through the training data). We observe two key trends. First, as training progresses, BLaST requires fewer FLOP to process the entire training set. Second, BLaST achieves higher accuracy with fewer FLOP, resulting in a superior accuracy-to-FLOP ratio compared to the dense network.
Figure 9. Training BLaST-90%/128x128 for ViT-B/16 on CIFAR-10. BLaST requires fewer PFLOP to process the entire training set as we sparsify the network. BLaST achieves higher accuracy with fewer FLOP, resulting in a better accuracy-to-FLOP ratio.
The previous results demonstrate that BLaST can be used to speed up the pretraining of Transformer based architectures with negligible loss of accuracy.
5.4. Ablation Studies
5.4.1. Use of Block Sparsity
Table 4 shows the perplexity of the GPT2 models for different block sizes $b\times b$ (see Section 3.2). Every measurement is performed by capping the maximum sparsity $s_{max}$ of the dense matrices of the MLP layers at 70%. Increasing the value of $s_{max}$ obscures the effect of varying block sizes due to a deterioration in accuracy (Hoefler2021, ). The sparsity mask is updated at every iteration (i.e. $step\_size$ from Listing LABEL:lst:loop-prune-and-grow is set to 1).
Model Dense
BLaST
1x1
BLaST
16x16
BLaST
32x32
BLaST
64x64
BLaST
128x128
GPT2-small 6.48 9.30 7.53 7.51 7.49 7.44
GPT2-medium 5.54 7.26 6.04 6.02 6.00 6.04
GPT2-large 5.02 6.43 5.47 5.42 5.49 5.45
GPT2-XL 4.79 5.95 5.15 5.11 5.13 5.13
Table 4. Pretraining perplexity of the test data set for various block sizes compared to dense with maximum sparsity $s_{max}$ at 70% for all BLaST measurements.
All models exhibit improved perplexity with blocked sparsity. This improvement can be attributed to the ”Difference” step shown in Figure 2. During the regrowth step in Figure 2, the regrown blocks alter the final sparsity of the new weight matrix $W_{new}$.
Figure 10 presents the ratio of regrown blocks to the total number of blocks for block sizes $b$ ranging from $1$ to $128$, based on GPT2-small measurements from Table 4. When $b=1$, the ratio attains its maximum value and exhibits the greatest variation. In contrast, the blocked variants with $b$ values from $16$ to $128$ show more stable and lower ratios, which correlates with improved perplexity.
The proportion of regrown blocks serves as an indicator of whether the pruning of the weight matrices introduces significant discontinuities in the optimization function (Hu2024, ). A lower proportion of regrown blocks suggests that the weight pruning is more consistent with the gradient’s optimization direction.
Figure 10. Proportion of regrown blocks for $b\in\{1,16,32,64,128\}$ on GPT2-small.
5.4.2. Sparsity Step Size
The $step\_size$ variable in Listing LABEL:lst:loop-prune-and-grow controls the number of iterations after which the sparse masks are generated using the blocked prune and grow algorithm from Section 3.2. The pruning function $S()$ in Figure 2 introduces a discontinuity in the optimization of the neural network (Hu2024a, ). The larger the interval between the mask generation steps, the more discontinuous the pruning function (Hu2024a, ). However, generating sparse masks introduces a non-trivial overhead into the overall computation. Table 5.4.2 tests the perplexity for $step\_size$ varying from 1 to 1,000.
{tblr}
colspec = c *8c, booktabs, cells = c $step\_size$ 1 2 5 10 25 50 100 1000
GPT2-small 7.51 7.52 7.51 7.50 7.53 7.54 7.50 9.83
GPT2-medium 6.02 6.02 6.03 6.03 6.03 6.01 6.03 6.51
GPT2-large 5.42 5.43 5.42 5.42 5.42 5.41 5.42 5.70
GPT2-XL 5.12 5.11 5.11 5.11 5.12 5.11 5.12 5.26
Table 5. Pre-training perplexity of BLaST versus pruning interval ($step\_size$) with fixed $32\times 32$ blocks. Results show robustness up to $step\_size=100$.
The perplexity is almost the same for a given model for all values of $step\_size$ except at 1,000 iterations. This experimentally proves that the overhead of generating sparse masks can be reduced without losing accuracy by using larger values of $step\_size$.
5.4.3. Changing the Sparsity Decay
Modifying $d$ in Equation 2 triggers the use of BLaST BSpMM routine earlier in training, speeding up the entire pretraining process. Table 6 illustrates the impact of varying $d$ for the GPT2-large model.
{tblr}
colspec = *6c, booktabs, cells = c Decay parameter $d$ 0 1000 4000 7000 9000
Perplexity 5.42 5.44 5.44 5.44 5.43
Table 6. BLaST perplexity for GPT2-large when changing $d$ from Equation 2 after training with 4.9 billion tokens.
The models in Table 6 are trained on 4.9 billion tokens, which corresponds to $m=10,000$. For the GPT2 models, the block size $b$ is set to $32\times 32$. Note that the drop in accuracy is negligible compared to the case of $d=0$, as shown in Table 4. This indicates that the BLaST BSpMM can be applied earlier in the pretraining process without sacrificing accuracy compared to $d=0$.
5.4.4. Effect of Dense Layers
Previous work (nvidia2025prune, ) has shown that pruning the first and the last dense layers has the greatest impact on the model’s performance. In Figure 11, we retain both the first and the last dense layers while applying sparsification to either the leftmost or the rightmost MLP blocks of the GPT2-small model. These results empirically indicate that the improvement in perplexity is greater when the dense MLP blocks are positioned on the right. Consequently, we employ dense layers on the right side of the model to enhance accuracy in the performance runs described in Section 5.3.2. The number of dense layers on the right side is denoted by $L$ in Table 2.
Figure 11. Impact on the perplexity when keeping a certain number of dense MLP blocks on the left, and right side. This ablation study determines the $L$ hyperparameter for pretraining in Section 5.3.
6. Related Work
Efficient ML
Various innovations, both algorithmic and hardware-based, have been proposed to reduce the time and cost associated with training and inference in large neural networks. These methods encompass quantization (nagel2021white, ; ashkboos2024quarot, ; dettmers2023spqr, ; frantar2023gptqaccurateposttrainingquantization, ; ashkboos2023quik, ), efficient dataset selection (okanovicrepeated, ; bartoldson2023compute, ; mirzasoleiman2020coresetsdataefficienttrainingmachine, ; yang2023towards, ; killamsetty2021grad, ; killamsetty2021glister, ), and weight pruning (Hoefler2021, ). In addition, tailored AI chips and GPUs with mixed precision compute units (Sean2023, ; Jouppi2017, ; Luo2024, ; Lee2024a, ; Abts2022, ; Markidis2018, ; Schieffer2024, ; Kalamkar2019, ; li2022efficient, ; Okanovic2024, ) have been designed to efficiently exploit memory bandwidth and parallelism, thereby accelerating matrix multiplications.
Dynamic low-rank approximation (DLRA) techniques (Koch2007, ; Schotthofer2022, ; Ceruti2022, ; Ceruti2022a, ) compress CNNs by leveraging rapidly decaying singular values of weight matrices. However, in transformer-based models these singular values decay slowly (Schotthofer2024, ; Cheng2005, ), reducing DLRA’s applicability. In contrast, weight pruning aims to remove individual weights, and has been successfully applied in both pre-training and post-training settings (Hu2024, ; Hubara2021, ; Zhou2021, ; Zhou2021a, ; Evci2021, ; Frankle2019, ; Chen2020, ; Child2019, ; Pan2024, ; Frantar2023, ; Ashkboos2024, ; Ivanov2023, ; Hu2021, ; Buyukakyuz2024, ; Chen2024, ). Structured sparsity approaches, such as 2:4 sparsity (Hu2024a, ; Mishra2021, ), coexist with unstructured methods based on the lottery ticket hypothesis (Frankle2019, ; Chen2021, ; You2022, ). Additional strategies include the use of straight-through estimators (Hu2024a, ) and transposable masks with minimum variance estimators (Hubara2021, ; Chmiel2024, ). Moreover, dynamic ’prune and grow’ methods that exploit first-order gradient momentum (Evci2021, ; Abdelfattah2024, ; cusparse, ; hipsparse, ) and knowledge distillation from large teacher networks (gou2021knowledge, ; hinton2015distillingknowledgeneuralnetwork, ; muralidharan2024compact, ) further enhance sparsity.
Matrix Multiplication
Sparse matrix multiplication on parallel architectures remains an active research area. Surveys summarize many effective techniques (gao2023, ; FilipponeReview, ), while recent work exploits dense MMA units to accelerate sparse computations (lu2023, ; niu2020, ; Zachariadis2020, ; Okanovic2024, ). For instance, DASP (lu2023, ) employs MMA units for SpMV acceleration, further boosted by CluB, which exposes dense substructures in sparse matrices. Reordering methods are key to improving data locality, bandwidth, and load balancing in sparse kernels (peng2020, ; Pichel2005, ; Pinar99, ; spMMHong, ; gianinazzi2024arrow, ). Graph-partitioning has boosted SpMV on multicore and GPU platforms (Trotter2023, ; Pichel2012, ; pichon2017, ; silva2017influence, ) using techniques such as Reverse Cuthill–McKee (RMC, ), approximate minimum degree (Amestoy2004, ), nested dissection (karypis1998fast, ), and Gray code ordering (Zhao2020, ). Recent advances, including community-based and just-in-time parallel reordering (Rabbit Order) (Balaji2023, ; rabbit, ) and others (gleinig2022optimal, ), further enhance cache locality and performance. Similarity-based reordering has been integrated into novel SpMM kernel designs. Hierarchical clustering via locality-sensitive hashing supports adaptive sparse tiling (peng2020, ; spMMHong, ), while similarity-based clustering improves both variable-size block multiplication (sylos2022, ), and the use of matrix units in block-based SpMM kernels (Okanovic2024, ).
7. Conclusion
This paper introduces BLaST, a general, scalable, and widely applicable end-to-end method that greatly reduces memory footprint, hardware cost of running contemporary extreme-scale LLMs, while simultaneously reducing runtime in all scenarios, both training- and inference-related. BLaST iteratively sparsifies the weights of the MLP layers inside the Transformer blocks, achieving up to 95% sparsity. Additionally, we introduce a highly efficient SpMM kernel that can be used as a standalone sparse linear algebra primitive, beyond ML workloads. We perform a comprehensive performance study of the proposed sparse kernel on state-of-the-art neural network architectures. We show a speedup of matrix multiplication of up to 16.7x, and we also demonstrate a speedup in the MLP block, achieving up to 8.8x improvement on the Llama-3.1 405B model. For Llama-3.2 1B, we show a 1.6x speedup in inference compared to the dense model, as well as a 1.1x speedup in pretraining. BLaST leads to memory savings of up to 3.12x during inference, translating into 2.9x lesser investment on GPUs and lower operating costs for data centers. We evaluate BLaST on 9 Transformer models across 8 datasets in both the text and vision domains, showing negligible accuracy losses compared to dense models. The presented results confirm that BLaST can be used as a drop-in method to scale up, and improve the efficiency of any Transformer-based architecture.